# The Universal Weight Subspace Hypothesis

### v1

@@ A.1 Related Work
Several lines of prior research support the core intuition behind our universal subspace hypothesis, though they do not provide a unified, scalable framework for identifying and leveraging such subspaces across architectures, tasks, and modalities. The Neural Tangent Kernel framework reinforces this idea, demonstrating that, in the infinite-width regime, training dynamics are governed by a kernel largely invariant to task specifics, implying the presence of common functional subspaces.  (jacot2018ntk). This result implies that training is implicitly constrained to a shared function space, suggesting the existence of low-dimensional structures that generalize across tasks. Complementing this, works in mechanistic interpretability has uncovered modular and recurring patterns that consistently re-emerge in independently trained models (olah2020zoom; chughtai2023toymodeluniversalityreverse), supporting the notion of structural universality in network representations. The closest related work is our previous work kaushik2025eigenloraxrecyclingadaptersprincipal, which investigates principal subspaces in low rank adapters for highly parameter efficient finetuning. That work considered a derivation of the analysis we present in this work. While works such as martin2025setolsemiempiricaltheorydeep; Mao_2024 have explored neural properties to explain generalization or loss landscape convergence, they do not address the convergence of parametric properties across distinct models trained on disjoint data. Importantly, in contrast to the abstract, weak, or speculative notions of universality presented in earlier works like chughtai2023toymodeluniversalityreverse; olah2020zoom - which focus primarily on representations and are arguably easier to demonstrate due to direct data dependency - ours is the first work to provide concrete evidence and a clear universal hypothesis at the neural parameter level.
Empirical studies further strengthen this perspective. The lottery ticket hypothesis (frankle2019lottery) demonstrates that overparameterized networks contain sparse subnetworks capable of matching full-model performance, implying that task-relevant information resides in a small, structured subset of weights. Similarly, mode connectivity studies (garipov2018mode) reveal that seemingly isolated optima in parameter space are often connected by low-loss paths, suggesting that task solutions lie on a shared manifold. In convolutional models, Krizhevsky et al. (Krizhevsky2012ImageNet) famously observed that early layers consistently learn Gabor-like filters, indicating a universal inductive bias in early representations. More recent works (guth24; guth2024universalityneuralencodingscnns) extends this observation to deeper layers, showing that certain eigenvectors of trained convolutional layers recur across networks trained on different datasets.
@@ A.2 Theoretical Analysis
Then from (minsker2017extensionsbernsteinsinequalityselfadjoint; koltchinskii2014concentrationinequalitiesmomentbounds) there exist absolute constants $C,c>0$ such that for every $\delta\in(0,1)$,
@@ Appendix B Universal Subspace Analysis
Similar methodology is followed for subspace analysis for both LoRA and classical weight models. In fact, LoRA analysis’ results can be theoretically extended to classical weights, as LoRA weights can be construed to be simple translations from a mean weight matrix. However, in order to solidify our universal subspace hypothesis, we conduct extensive experiments for both types of models. LoRA is chosen because of the recent spurt in the availability of LoRA models trained on diverse kinds of datasets and models. We do this universal subspace analysis on all weight parameters in every neural network layer except the first (or few initial) and last neural network layer. This is because these layers may differ across models due to differences in input shapes and types, loss functions, and the tasks being trained. We also focus our analysis on linear/fully-connected and matrix weights, as the analysis done on these are straightforward and the results observed can be trivially extended to other types of neural parameters (ma2017equivalencefullyconnectedlayer).
@@
refers to the residual subspace that remains after removing the top $k$ principal directions associated with the low-rank universal subspace. This subspace is orthogonal to the universal subspace and serves as a control for evaluating the uniqueness and effectiveness of the learned shared subspace. To make computation tractable when the residual subspace is high-dimensional, we focus on the top components beyond rank $k$, as computing a full SVD is often impractical. This approximation is justified, since the lower components typically capture noise, which has been shown to degrade performance (sharma_laser_2023).
@@
As shown in all eigenvalue (scree) plots, a trivial way to choose is a simple visual inspection, since we can see a discontinuity in the spectral analysis. Another way is to define a threshold on the explained variance, all components whose explained variance is close to zero <.01 are considered secondary subspace, and can be discarded. A more structured way is to define an optimal singular value threshold for the HOSVD, as found by previous works (gavish2014optimalhardthresholdsingular).
@@ B.1 Lower rank shared universal subspaces within low rank adaptation (LoRA) models
In our first experimental analysis, we use 500 LoRA models trained on distinct Natural Instructions (wang-etal-2022-super) using Mistral-7B-Instruct-v0.2 (jiang2023mistral7b) as the base (brüelgabrielsson2024compressserveservingthousands). Please refer to  brüelgabrielsson2024compressserveservingthousands for more details on how the LoRA models were trained.
@@
Our second experiment involves the complex and multimodal task of Text-to-Image generation using the Stable Diffusion-XL model sdxl. We extract our low rank universal subspace from publicly available LoRA models on HuggingFace repository von-platen-etal-2022-diffusers - LABEL:tab:diffusion_data lists all the SDXL models that we used to extract the Universal Subspace. As can be seen in LABEL:tab:diffusion_data, the models range wildly in styles on which they were finetuned. The fact that all these diverse models can be represented by a single low rank universal subspace model strongly verifies our hypothesis. We use top 16 components and 30 denoising steps. For each experiment model shown in Table 1 and Figure 5, that LoRA model is reconstructed using a universal subspace created using rest of the available LoRA adapters, essentially confirming the generalization capability of this subspace.
We then use this single SDXL universal subspace to generate images with similar styles to evaluate whether this subspace is capable of doing so, by projecting randomly chosen LoRA models into this subspace. Figure 5 shows that our universal subspace matches the visual quality and style nuances of individual LoRAs, resulting in significant memory savings. Table 1 shows quantitative results for our Universal subspace in terms of CLIP scores, where interestingly we can see that our Universal Subspace outperforms the individual LoRA models. This improvement may be attributed to our Universal SDXL removing noise from the subspace - a phenomenon previously observed by sharma_laser_2023. The styles used in Table˜1, which are in LABEL:tab:diffusion_data are (from Style 1 to Style 10) Ukiyo-e Style, Todd Hildo Style , Olly Moss Style , Needlepoint Style , Studio Ghibli Style, Surreal Harmony Style , Dressed Animal Style , Lascaux Cave Art Style , Kirigami Style , Yaacov Agam Style.
@@ B.2 Low rank shared universal subspaces in classical weights
In order to further solidify the evidence for our universal subspace hypothesis, we show that this universality does extend beyond adapter models to conventional weights. We do not focus on convolutional weight parameters as they can simply be equated with fully connected layers (ma2017equivalencefullyconnectedlayer), and have been shown, in limited scope, to match Gabor-like filters (Krizhevsky2012ImageNet). Therefore, our analysis trivially extends to these kinds of parameters as well. However, there are a few practical differences between the low rank adapter and classical weight subspace analysis. The classical weight subspace analysis is more computationally expensive relative to the LoRA one due to high dimensionality of the parameters, but in effect, same. Additionally, the number of sufficiently well trained models is understandably fewer than LoRA models. Further, there is also higher variance in terms of model quality in the classical weights as it is harder to optimize these models as compared to LoRA which often are optimized from a good initialization point (the pretrained base model). An outcome of this is that the universal subspace approximation that we obtain from the publicly available pretrained models are noisier than their LoRA counterparts. Inspite of this, our universal subspace hypothesis remains validated.
To further support our universal subspace hypothesis, we extend our analysis beyond adapter models to standard full-rank weights. We exclude convolutional parameters from explicit consideration, as they are functionally equivalent to fully connected layers under certain conditions (ma2017equivalencefullyconnectedlayer), and their learned representations (e.g., Gabor-like filters) have been studied, in limited scope, in prior work (Krizhevsky2012ImageNet). Consequently, our analysis generalizes naturally to convolutional weights as well.
@@
We collected Flan-T5 models fine-tuned on individual datasets from the GLUE (glue) benchmark. We extract the joint subspace from these models and trends similar to those observed above are seen. This shows that across diverse datasets and tasks a low-rank subspace emerges.
@@ Appendix C Finding universal subspaces and applying them to future tasks
In this section, we present two tasks, GLUE (glue) and Image Classification. For each experiment, the joint subspace is created using all other models in subset. For Image Classification, we use $k=4$ and train only 8 epochs using learning rate of 1e-4. Importantly, only the coefficients are trained for the experiment. It is important to note that our shared subspace model performs quite well despite using very few (4-5) models to extract the subspace. For GLUE, we use 16-32 components for our subspace, with learning rate of 4e-4, batch size of 64, and 30-80 epochs for each task. In addition, it is likely that our model might perform similarly or better if trained longer or with optimized hyperparameters.

### v2

@@ A.1 Related Work
Several lines of prior research support the core intuition behind our universal subspace hypothesis, though they do not provide a unified, scalable framework for identifying and leveraging such subspaces across architectures, tasks, and modalities. The Neural Tangent Kernel framework reinforces this idea, demonstrating that, in the infinite-width regime, training dynamics are governed by a kernel largely invariant to task specifics, implying the presence of common functional subspaces.  (Jacot et al., 2018). This result implies that training is implicitly constrained to a shared function space, suggesting the existence of low-dimensional structures that generalize across tasks. Complementing this, works in mechanistic interpretability has uncovered modular and recurring patterns that consistently re-emerge in independently trained models (Olah et al., 2020; Chughtai et al., 2023), supporting the notion of structural universality in network representations. The closest related work is our previous work Kaushik et al. (2025), which investigates principal subspaces in low rank adapters for highly parameter efficient finetuning. That work considered a derivation of the analysis we present in this work. While works such as Martin & Hinrichs (2025); Mao et al. (2024) have explored neural properties to explain generalization or loss landscape convergence, they do not address the convergence of parametric properties across distinct models trained on disjoint data. Importantly, in contrast to the abstract, weak, or speculative notions of universality presented in earlier works like Chughtai et al. (2023); Olah et al. (2020) - which focus primarily on representations and are arguably easier to demonstrate due to direct data dependency - ours is the first work to provide concrete evidence and a clear universal hypothesis at the neural parameter level.
Empirical studies further strengthen this perspective. The lottery ticket hypothesis (Frankle & Carbin, 2019) demonstrates that overparameterized networks contain sparse subnetworks capable of matching full-model performance, implying that task-relevant information resides in a small, structured subset of weights. Similarly, mode connectivity studies (Garipov et al., 2018) reveal that seemingly isolated optima in parameter space are often connected by low-loss paths, suggesting that task solutions lie on a shared manifold. In convolutional models, Krizhevsky et al. (Krizhevsky et al., 2012) famously observed that early layers consistently learn Gabor-like filters, indicating a universal inductive bias in early representations. More recent works (Guth et al., 2024; Guth & Ménard, 2024) extends this observation to deeper layers, showing that certain eigenvectors of trained convolutional layers recur across networks trained on different datasets.
@@ A.2 Theoretical Analysis
Then from (Minsker, 2017; Koltchinskii & Lounici, 2014) there exist absolute constants $C,c>0$ such that for every $\delta\in(0,1)$,
@@ Appendix B Universal Subspace Analysis
Similar methodology is followed for subspace analysis for both LoRA and classical weight models. In fact, LoRA analysis’ results can be theoretically extended to classical weights, as LoRA weights can be construed to be simple translations from a mean weight matrix. However, in order to solidify our universal subspace hypothesis, we conduct extensive experiments for both types of models. LoRA is chosen because of the recent spurt in the availability of LoRA models trained on diverse kinds of datasets and models. We do this universal subspace analysis on all weight parameters in every neural network layer except the first (or few initial) and last neural network layer. This is because these layers may differ across models due to differences in input shapes and types, loss functions, and the tasks being trained. We also focus our analysis on linear/fully-connected and matrix weights, as the analysis done on these are straightforward and the results observed can be trivially extended to other types of neural parameters (Ma & Lu, 2017).
@@
refers to the residual subspace that remains after removing the top $k$ principal directions associated with the low-rank universal subspace. This subspace is orthogonal to the universal subspace and serves as a control for evaluating the uniqueness and effectiveness of the learned shared subspace. To make computation tractable when the residual subspace is high-dimensional, we focus on the top components beyond rank $k$, as computing a full SVD is often impractical. This approximation is justified, since the lower components typically capture noise, which has been shown to degrade performance (Sharma et al., 2023).
@@
As shown in all eigenvalue (scree) plots, a trivial way to choose is a simple visual inspection, since we can see a discontinuity in the spectral analysis. Another way is to define a threshold on the explained variance, all components whose explained variance is close to zero <.01 are considered secondary subspace, and can be discarded. A more structured way is to define an optimal singular value threshold for the HOSVD, as found by previous works (Gavish & Donoho, 2014).
@@ B.1 Lower rank shared universal subspaces within low rank adaptation (LoRA) models
In our first experimental analysis, we use 500 LoRA models trained on distinct Natural Instructions (Wang et al., 2022) using Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as the base (Brüel-Gabrielsson et al., 2024). Please refer to  Brüel-Gabrielsson et al. (2024) for more details on how the LoRA models were trained.
@@
Our second experiment involves the complex and multimodal task of Text-to-Image generation using the Stable Diffusion-XL model Podell et al. (2023). We extract our low rank universal subspace from publicly available LoRA models on HuggingFace repository von Platen et al. (2022) - LABEL:tab:diffusion_data lists all the SDXL models that we used to extract the Universal Subspace. As can be seen in LABEL:tab:diffusion_data, the models range wildly in styles on which they were finetuned. The fact that all these diverse models can be represented by a single low rank universal subspace model strongly verifies our hypothesis. We use top 16 components and 30 denoising steps. For each experiment model shown in Table 1 and Figure 5, that LoRA model is reconstructed using a universal subspace created using rest of the available LoRA adapters, essentially confirming the generalization capability of this subspace.
We then use this single SDXL universal subspace to generate images with similar styles to evaluate whether this subspace is capable of doing so, by projecting randomly chosen LoRA models into this subspace. Figure 5 shows that our universal subspace matches the visual quality and style nuances of individual LoRAs, resulting in significant memory savings. Table 1 shows quantitative results for our Universal subspace in terms of CLIP scores, where interestingly we can see that our Universal Subspace outperforms the individual LoRA models. This improvement may be attributed to our Universal SDXL removing noise from the subspace - a phenomenon previously observed by Sharma et al. (2023). The styles used in Table˜1, which are in LABEL:tab:diffusion_data are (from Style 1 to Style 10) Ukiyo-e Style, Todd Hildo Style , Olly Moss Style , Needlepoint Style , Studio Ghibli Style, Surreal Harmony Style , Dressed Animal Style , Lascaux Cave Art Style , Kirigami Style , Yaacov Agam Style.
@@ B.2 Low rank shared universal subspaces in classical weights
In order to further solidify the evidence for our universal subspace hypothesis, we show that this universality does extend beyond adapter models to conventional weights. We do not focus on convolutional weight parameters as they can simply be equated with fully connected layers (Ma & Lu, 2017), and have been shown, in limited scope, to match Gabor-like filters (Krizhevsky et al., 2012). Therefore, our analysis trivially extends to these kinds of parameters as well. However, there are a few practical differences between the low rank adapter and classical weight subspace analysis. The classical weight subspace analysis is more computationally expensive relative to the LoRA one due to high dimensionality of the parameters, but in effect, same. Additionally, the number of sufficiently well trained models is understandably fewer than LoRA models. Further, there is also higher variance in terms of model quality in the classical weights as it is harder to optimize these models as compared to LoRA which often are optimized from a good initialization point (the pretrained base model). An outcome of this is that the universal subspace approximation that we obtain from the publicly available pretrained models are noisier than their LoRA counterparts. Inspite of this, our universal subspace hypothesis remains validated.
To further support our universal subspace hypothesis, we extend our analysis beyond adapter models to standard full-rank weights. We exclude convolutional parameters from explicit consideration, as they are functionally equivalent to fully connected layers under certain conditions (Ma & Lu, 2017), and their learned representations (e.g., Gabor-like filters) have been studied, in limited scope, in prior work (Krizhevsky et al., 2012). Consequently, our analysis generalizes naturally to convolutional weights as well.
@@
We collected Flan-T5 models fine-tuned on individual datasets from the GLUE (Wang et al., 2019) benchmark. We extract the joint subspace from these models and trends similar to those observed above are seen. This shows that across diverse datasets and tasks a low-rank subspace emerges.
@@ Appendix C Finding universal subspaces and applying them to future tasks
In this section, we present two tasks, GLUE (Wang et al., 2019) and Image Classification. For each experiment, the joint subspace is created using all other models in subset. For Image Classification, we use $k=4$ and train only 8 epochs using learning rate of 1e-4. Importantly, only the coefficients are trained for the experiment. It is important to note that our shared subspace model performs quite well despite using very few (4-5) models to extract the subspace. For GLUE, we use 16-32 components for our subspace, with learning rate of 4e-4, batch size of 64, and 30-80 epochs for each task. In addition, it is likely that our model might perform similarly or better if trained longer or with optimized hyperparameters.